The math contest called the Fermat is out of 150 points. The mean score this year is 91 points with a standard deviation of 12. Students in the top 25% receive a medal. Assuming the scores are normally distributed, what mark is required to receive a medal? *




1
Expert's answer
2021-07-26T14:26:08-0400


Given that:

"\mu=91"

"\sigma=12"

"n=150"


"\lambda=?"


"Area: 1-0.25=0.75"

The "Z-Score" of "0.75" is "0.67".

"Z-score=\frac{\lambda -\mu }{\frac{\sigma }{\sqrt{n}}}"

"0.67=\frac{\lambda -91}{\frac{12}{\sqrt{150}}}"

"\lambda -91=0.67\times \frac{12}{\sqrt{150}}"

"\lambda =\left(0.67\times \frac{12}{\sqrt{150}}\right)+91"

"\lambda =91.6565\approx 91.66"


Therefore, "91.66" are required to receive a medal.